# REArtGS: Reconstructing and Generating Articulated Objects via 3D Gaussian Splatting with Geometric and Motion Constraints

### v2

@@ Abstract
Articulated objects, as prevalent entities in human life, their 3D representations play crucial roles across various applications. However, achieving both high-fidelity textured surface reconstruction and dynamic generation for articulated objects remains challenging for existing methods. In this paper, we present REArtGS, a novel framework that introduces additional geometric and motion constraints to 3D Gaussian primitives, enabling high-quality textured surface reconstruction and generation for articulated objects. Specifically, given multi-view RGB images of arbitrary two states of articulated objects, we first introduce an unbiased Signed Distance Field (SDF) guidance to regularize Gaussian opacity fields, enhancing geometry constraints and improving surface reconstruction quality. Then we establish deformable fields for 3D Gaussians constrained by the kinematic structures of articulated objects, achieving unsupervised generation of surface meshes in unseen states. Extensive experiments on both synthetic and real datasets demonstrate our approach achieves high-quality textured surface reconstruction for given states, and enables high-fidelity surface generation for unseen states. Codes will be released within the next four months and the project website is at https://sites.google.com/view/reartgs/home.
@@ 3.3 Optimization and Textured Mesh Extraction
Table 1: Quantitative results for the surface reconstruction quality on PartNet-Mobility dataset. We bold the best results and underline the second best results.
@@
Table 2: Quantitative results for the surface generation quality on PartNet-Mobility dataset. We bold the best results and underline the second best results.
@@ 4.3 Mesh Generation Performance
We report the quantitative results of mesh reconstruction in Table. 2 and present the qualitative results Fig. 5. Please refer to our supplementary material for more qualitative results. Our REArtGS achieves the best mean results across all metrics with 5.41, 1.063, 0.276 and 0.670. Compared to PARIS, our method exhibits smoother surface generation results, as shown in Fig. 5. It is mainly because PARIS lacks a reasonable geometry initiation and directly utilizes the composite rendering, leading to the spatial overlap between the neural radiation fields of the start and end states. In contrast, our approach utilizes the accurate dense points from the reconstruction stage and optimizes the deformable fields to yield dynamic results, maintaining high-quality surface during generation.
@@ 4.5 Generalization to the Real World
Box PARIS 4.69 1.651 0.097 0.843 3.98 2.131 0.306 1.055
REArtGS 2.49 0.898 0.205 0.671 1.28 1.425 0.559 0.845
Stapler PARIS 0.23 0.104 0.551 0.165 0.47 0.198 0.540 0.194
REArtGS 0.24 0.037 0.728 0.137 0.34 0.044 0.582 0.151
Scissor PARIS 0.18 0.302 0.710 0.314 0.18 0.329 0.424 0.310
REArtGS 0.08 0.017 0.902 0.093 0.08 0.003 0.899 0.042
Cutter PARIS 2.56 1.371 0.570 0.302 45.02 1.684 0.309 1.183
REArtGS 0.13 0.016 0.874 0.090 51.56 1.559 0.228 0.884
Drawer PARIS 30.63 4.568 0.062 1.577 33.94 3.865 0.091 1.430
REArtGS 10.73 2.607 0.138 1.142 11.71 1.739 0.209 0.933
Mean PARIS 7.66 1.599 0.398 0.640 16.72 1.641 0.334 0.591
REArtGS 3.42 0.715 0.569 0.427 12.99 0.954 0.495 0.571
Table 5: The quantitative results on real-world AKB-48 dataset.

### v3

@@ Abstract
Articulated objects, as prevalent entities in human life, their 3D representations play crucial roles across various applications. However, achieving both high-fidelity textured surface reconstruction and dynamic generation for articulated objects remains challenging for existing methods. In this paper, we present REArtGS, a novel framework that introduces additional geometric and motion constraints to 3D Gaussian primitives, enabling high-quality textured surface reconstruction and generation for articulated objects. Specifically, given multi-view RGB images of arbitrary two states of articulated objects, we first introduce an unbiased Signed Distance Field (SDF) guidance to regularize Gaussian opacity fields, enhancing geometry constraints and improving surface reconstruction quality. Then we establish deformable fields for 3D Gaussians constrained by the kinematic structures of articulated objects, achieving unsupervised generation of surface meshes in unseen states. Extensive experiments on both synthetic and real datasets demonstrate our approach achieves high-quality textured surface reconstruction for given states, and enables high-fidelity surface generation for unseen states. Codes will be released after acceptance and the project website is at https://sites.google.com/view/reartgs/home.
@@ 3.3 Optimization and Textured Mesh Extraction
Table 1: Quantitative results for the surface reconstruction quality on PartNet-Mobility dataset. We bold the best results and underline the second best results. * means we implement it without depth supervision for fair comparison.
@@
Table 2: Quantitative results for the surface generation quality on PartNet-Mobility dataset. We bold the best results and underline the second best results. * means we implement it without depth supervision for fair comparison.
@@ 4.3 Mesh Generation Performance
We report the quantitative results of mesh reconstruction in Table. 2 and present the qualitative results Fig. 5. Please refer to our supplementary material for more qualitative results. Our REArtGS achieves the best mean results across all metrics with 4.10, 1.063, 0.298 and 0.670. Compared to PARIS, our method exhibits smoother surface generation results, as shown in Fig. 5. It is mainly because PARIS lacks a reasonable geometry initiation and directly utilizes the composite rendering, leading to the spatial overlap between the neural radiation fields of the start and end states. In contrast, our approach utilizes the accurate dense points from the reconstruction stage and optimizes the deformable fields to yield dynamic results, maintaining high-quality surface during generation.
@@ 4.5 Generalization to the Real World
Box PARIS [15] 4.69 1.651 0.097 0.843 3.98 2.131 0.306 1.055
ArtGS [18] 8.61 0.967 0.143 0.697 10.23 1.606 0.309 0.851
REArtGS (Ours) 2.49 0.898 0.205 0.671 1.28 1.425 0.559 0.845
Stapler PARIS [15] 0.23 0.104 0.551 0.165 0.47 0.198 0.540 0.194
ArtGS [18] 51.04 1.364 0.476 0.827 5.13 0.527 0.533 0.514
REArtGS (Ours) 0.24 0.037 0.728 0.137 0.34 0.044 0.582 0.151
Scissor PARIS [15] 0.18 0.302 0.710 0.314 0.18 0.329 0.424 0.310
ArtGS [18] 11.43 0.073 0.588 0.192 5.53 0.089 0.600 0.213
REArtGS (Ours) 0.08 0.017 0.902 0.093 0.08 0.003 0.899 0.042
Cutter PARIS [15] 2.56 1.371 0.570 0.302 45.02 1.684 0.309 1.183
ArtGS [18] 5.65 0.453 0.644 0.477 65.64 1.298 0.093 0.806
REArtGS (Ours) 0.13 0.016 0.874 0.090 51.56 1.559 0.228 0.884
Drawer PARIS [15] 30.63 4.568 0.062 1.577 33.94 3.865 0.091 1.430
ArtGS [18] 19.41 7.478 0.086 1.935 12.60 1.624 0.094 0.911
REArtGS (Ours) 10.73 2.607 0.138 1.142 11.71 1.739 0.209 0.933
Eyeglasses PARIS [15] 0.16 1.582 0.637 0.212 0.18 1.550 0.609 1.284
ArtGS [18] 27.49 1.384 0.557 0.832 12.12 1.592 0.747 0.955
REArtGS (Ours) 0.04 0.018 0.969 0.096 0.04 1.568 0.974 0.886
Mean PARIS [15] 6.41 1.336 0.438 0.569 13.96 1.626 0.380 0.707
ArtGS [18] 20.61 1.953 0.416 0.826 18.54 1.123 0.396 0.708
REArtGS (Ours) 2.86 0.599 0.636 0.372 10.83 1.056 0.575 0.624
Table 5: The quantitative results on real-world AKB-48 dataset. We bold the best results and underline the second best results. * means we implement it without depth supervision for fair comparison.